OpenStudy (anonymous):

The gas tank of a truck is a cylinder 4.5ft long with a diameter of 2.5ft . At the gas station, a pump pours gas at the rate of 5ft^3 per minute. How many minutes does it take to fill the empty tank with that pump? Use the value 3.14 for Pi , and round your answer to the nearest minute.

OpenStudy (dumbcow):

Volume of cylinder = (pi*r^2)h h = 4.5 r = 1.25 Volume = rate*time Volume = 5*time time = Volume/5

OpenStudy (anonymous):

r= radius

OpenStudy (dumbcow):

yes

OpenStudy (anonymous):

allright hold on a sec

OpenStudy (anonymous):

i got 35.325

OpenStudy (dumbcow):

for the Volume? should be around 22

OpenStudy (anonymous):

hmm ok

OpenStudy (dumbcow):

V = (3.14)(1.25^2)(4.5)